5r-3s=17
3r+5s=51
Elimination Method

Found 2 solutions by stanbon, mananth:
Answer by stanbon(75887)

5r-3s=17
3r+5s=51
-----
Multiply 1st by 3
Multiply 2nd by 5
-----
15r - 9s = 3*17
15r +25s = 5*51
------------
Subtract and solve for "s":
34s = 204
---
s = 6
----
Solve for "r":
5r - 3*6 = 17
5r = 17+18
5r = 35
r = 7
-------
Cheers,
Stan H.
-------------
=============

Answer by mananth(16949)

5 r -3 s = 17 .............1
3 r + 5 s = 51 .............2

multiply (1)by 5
Multiply (2) by 3
25 r + -15 s = 85
9 r + 15 s = 153

Add the two equations

34 r + 0 s = 238

/ 34

r = 7

plug value of r in (1)

5 r + -3 s = 17
35 + -3 s = 17
-3 s = 17-35
-3 s = -18
s = 6